# amenability and fixed points in cones

### v1

@@ Abstract.
We characterize the groups for which these generalized probabilities can be stationary, respectively invariant. This leads to a new criterion for amenability and for fixed points in cones.
@@ 1. Introduction
An elementary case is when we have a positive linear functional $p$ on $V$ that is faithful in the sense that $p(v)>0$ for all $v\gneqq 0$. Then $r(u,v)=p(u)/p(v)$ is a vector pricing based on the “gold standard” pp.
@@ 1.2. Motivation: conditional probability
In the unconditional case, there is a well-known identification between finitely additive probability measures and normalized positive integrals, called means, on the space $\ell^{\infty}$ of all bounded functions. Therefore we ask:
@@
These axioms suggest a source of conditional means: take a vector pricing $r$ on $V=\ell^{\infty}(X)$ and define $P(f|h)=r(f,h)$ when $0\leq f\leq h$. This is a functional generalization of the approach studied by Armstrong–Sudderth [AS89] in the case of probabilities. We will see that vector pricings and conditional means are essentially equivalent for $f\leq h$ (Section˜3.1).
Remarkably, conditional means are not simply an extension of conditional probabilities, unlike in the unconditional case. We begin with a warning pennant based on the recent construction of finitely additive stationary measures [AHP+25]. Recall that the random walk defined by a distribution $\mu$ on a (discrete) group $G$ is irreducible as a Markov chain iff $\mu$ is non-degenerate, i.e. its support generates $G$ as a semigroup. A conditional mean is $\mu$-stationary if
$P(\mu*f|h)=P(f|h)$
holds whenever defined.
@@
The non-amenable case of ˜C follows rapidly from considering a generalized Green function. The amenable case seems more difficult and motivated us to establish an “eigenfunctional theorem”, see [Mon25], which enters the proof.
Since amenability is equivalent to admitting a $G$-invariant mean, we now have to ask when $G$ admits a $G$-invariant conditional mean, i.e. $P(gf|h)=P(f|h)$ for all $g\in G$, where defined. The answer involves the fixed-point property for cones introduced in [Mon17]. For the time being, we only mention that a number of equivalent characterization are established in [Mon17], among which a translate property that goes back to Dixmier [Dix50], Geenleaf [Gre69] and Rosenblatt [Ros73]. It is further shown in [Mon17] that this fixed-point property is preserved under subgroups, quotients, directed unions, central extension and products with subexponential groups. It fails notably for groups containing free semigroups, e.g. the lamplighter $(\mathbf{Z}/2)\wr\mathbf{Z}$.
@@
This would be parallel to the difference between ˜B and ˜C if stationarity is replaced by invariance.
@@ 1.3. Invariance for pricings, means and functionals
There is no obvious cone to which we could apply the fixed-point property to get an invariant vector pricing; for instance, axiom (VP2) is not a convex condition. Therefore the main step for ˜F will be to reduce the question to numerous cones of linear functionals using the following result.
@@
Let $G$ be a group with a representation on an ordered vector space $V$. Suppose that for every non-zero positive vector $v\in V$, the $G$-invariant ideal $V_{Gv}$ generated by the orbit GvGv admits a non-zero $G$-invariant positive linear functional.
@@ 1.4. Domains of definition
The critical reader might question two choices that we made from the start. First, we defined conditional means $P(u|v)$ only when $u\leq v$, whereas the conditional probability $P(A|B)\in[0,1]$ is classically defined for any $A\subseteq X$, not only for $A\subseteq B$. In fact, Rényi’s original axioms [Rén55] are that $P(\cdot|B)$ is a probability measure on $X$ such that
@@
Moreover, $P(\boldsymbol{1}_{A}|\boldsymbol{1}_{B})=P(\boldsymbol{1}_{A\cap B}|\boldsymbol{1}_{B})$ holds for arbitrary $A,B\neq\varnothing$ when $V$ the space of step-functions.
@@
Section˜2 sets the table with notation and background. Section˜3 introduces partial positive functionals and relates them to vector pricings and conditional means. A vectorial analogue of the Rényi order on measures is also introduced. Section˜4 discusses invariance. Section˜5 deals with stationarity. Section˜6 squares away questions of domains of definition.
@@ 2.1. Order
We also note that the dual of a representation is strictly speaking a representation of the opposite semigroup on the dual space, or equivalently a right action. We shall however abuse notation and speak of dual $G$-actions; for groups, this is of course the same as the usual action by the inverse element.
@@
Let $A\subseteq V$ be any subset. The ideal generated by AA is the intersection $V_{A}$ of all ideals of $V$ containing AA. When $A=\{v\}$, the simpler notation $V_{v}$ is used. We record a few basic facts that will be used throughout.
@@
For (i), any ideal containing AA must contain $\operatorname{Span}(A)$ and hence also the right hand side. For the reverse inclusion, it suffices to observe that the right hand side is an ideal; it follows that it contains $V_{A}$.
@@
We call $V$ singular if there is $v\gneqq 0$ such that $J(v)=0$ for every positive linear functional $J$ on $V$. This does not happen in most familiar spaces, but there are extreme examples.
Example $2.$
@@
$2.$ The definition of vector pricings
@@ 4. Invariance
For the record, we clarify the definition of invariance for vector pricings. Although we are mainly working with groups, we state this more generally for semigroups because this will be convenient for handling stationarity.
@@ 4.2. Invariant pricings
Let $G$ be a semigroup with a representation on an ordered vector space $V$. Suppose that for every non-zero positive vector $v\in V$, the $G$-invariant ideal $V_{Gv}$ generated by the orbit GvGv admits a non-zero $G$-invariant positive linear functional.
@@
We define a collection $\mathscr{C}$ of partial positive functionals as follows. Given any non-zero $v\geq 0$, let $V_{Gv}$ be the ideal generated by the orbit GvGv; this is a $G$-invariant ideal. By assumption, there exists a non-zero $G$-invariant positive linear map $J\colon V_{Gv}\to\mathbf{R}$. We note that $J(v)\neq 0$; indeed otherwise by invariance $J(Gv)=\{0\}$ so that $J$ would vanish on the entire ideal $V_{Gv}$.
@@
Finally, the “local boundedness” condition postulates the existence of a non-zero element $x\in E^{+}$ whose orbit is bounded in the sense of topological vector spaces. Since the representation is order-bounded, there is $w\in V$ with $gv\leq w$ for all $g\in G$. Since the order is non-singular, we can find a positive linear functional $K\colon V\to\mathbf{R}$ with $K(v)\neq 0$. Now we claim that the restriction of $K$ to $V_{Gv}$ is the desired element $x\in E^{+}$. Indeed, weak-* boundedness of GxGx means that any basic neighbourhood of 0 of the form
@@
we must absorb GxGx, i.e. $Gx\subseteq tU$ must hold for all sufficiently large $t>0$. Given such $v_{i}$ and $\epsilon$, we can find for each $i$ elements $g_{i,j}\in G$ with $j=1,\ldots,k_{i}$ and $\pm v_{i}\leq\sum_{j=1}^{k_{i}}g_{i,j}v$ because $v_{i}\in V_{Gv}$ (again ˜2.1(ii)). Now
@@ 5. Stationarity
We denote its orbits by $\langle\mu\rangle v=\{\mu^{*n}*v:n\in\mathbf{N}\}$ and observe that the orbit ideal $V_{\langle\mu\rangle v}$ is contained in the $G$-invariant ideal $V_{Gv}$ whenever $v\in V^{+}$. In particular, if the $G$-representation is order-bounded, then $\mu$ has order-bounded orbits. Moreover, if $\mu$ is non-degenerate, then we have $V_{\langle\mu\rangle v}=V_{Gv}$ because every gvgv is in $V_{\mu^{*n}v}$ when $n$ is large enough relative to $g\in G$.
@@
In our current setting, this can be combined with Kersten’s criterion to obtain the following result.
@@ 6.1. Step-functions
Given a non-empty subset $A\subseteq X$, we consider $P(\cdot|A)$ as a finitely additive probability measure on AA, which has therefore a unique extension to a mean $P^{\prime}(\cdot|A)$ on the space $\ell^{\infty}(A)$. Thus in particular $P^{\prime}(\boldsymbol{1}_{B}|A)=P(B|A)$ when $B\subseteq A$.
@@ 6.2. Globalizing conditional means
$\forall u\in V^{+}:\kern 8.53581ptp_{v}(u)=\sup\{u\wedge nu:n\in\mathbf{N}\}\in V_{v}.$
@@
$P(p_{v}(u)|v)=P\big(\sup\{u\wedge nu:n\in\mathbf{N}\}\big|v\big)=\sup_{n\in\mathbf{N}}P\left(u\wedge nv|v\right),$
@@ 6.3. Staying positive
˜6.6 justifies the above claim because any natural extension of an invariant vector pricing provided by ˜E or ˜F for a non-trivial group would still be invariant and thus contradict ˜$6.$
@@
again absurd. All values of $r$ appearing in the above rather dim sum are 11, then 0, so that no hypothetical $\infty-\infty$ could be used as a fig leaf. ∎

### v2

@@ Abstract.
We characterize the groups for which these generalized probabilities can be stationary, respectively invariant. Our results deviate from the setting of classical probability; this leads to a new criterion for amenability and for fixed points in cones.
@@ 1.1. The price of vectors
An elementary case is when we have a positive linear functional $p$ on $V$ that is faithful in the sense that $p(v)>0$ for all $v\gneqq 0$. Then $r(u,v)=p(u)/p(v)$ is a vector pricing based on the “gold standard” $p$.
@@ 1.2. Motivation: conditional probability
In the unconditional case, there is a well-known identification between finitely additive probability measures and normalized positive integrals, called means, or states, on the space $\ell^{\infty}$ of all bounded functions. Therefore we ask:
@@
These axioms suggest an obvious source of conditional means: take a vector pricing $r$ on $V=\ell^{\infty}(X)$ and define $P(f|h)=r(f,h)$ when $0\leq f\leq h$. This is a functional generalization of the approach studied by Armstrong–Sudderth [AS89] in the case of probabilities. We will see that vector pricings on $\ell^{\infty}(X)$ and conditional means on $X$ are essentially equivalent (Section˜3.1).
Remarkably, conditional means are not simply an extension of conditional probabilities, unlike in the unconditional case. We begin with a warning pennant based on the recent construction of stationary measures [AHP+25]. Recall that the random walk defined by a distribution $\mu$ on a (discrete) group $G$ is irreducible as a Markov chain iff $\mu$ is non-degenerate, i.e. its support generates $G$ as a semigroup. A conditional probability or mean is $\mu$-stationary if
$P(\mu*\cdot|\cdot)=P(\cdot|\cdot)$
holds whenever defined, where * denotes convolution.
@@
The non-amenable case of ˜C follows rapidly from considering a generalized Green function. The amenable case is substantially more difficult and motivated us to establish an “eigenfunctional theorem”, see [Mon25], which enters the proof.
Since amenability is equivalent to admitting a $G$-invariant mean, we now must ask: when does $G$ admit a $G$-invariant conditional mean? i.e. $P(gf|h)=P(f|h)$ for all $g\in G$, where defined. The answer involves the fixed-point property for cones introduced in [Mon17]. For the time being, we only mention that a number of equivalent characterization are established in [Mon17], among which a translate property that goes back to Dixmier [Dix50], Geenleaf [Gre69] and Rosenblatt [Ros73]. It is further shown in [Mon17] that this fixed-point property is preserved under subgroups, quotients, directed unions, central extension and products with subexponential groups. It fails notably for groups containing free semigroups, e.g. the lamplighter $(\mathbf{Z}/2)\wr\mathbf{Z}$.
@@
This would be parallel to the difference between ˜B and ˜C if stationarity is replaced by invariance. We view this as evidence for the conjecture.
@@ 1.3. Invariance for pricings, means and functionals
There is no obvious cone to which we could apply the fixed-point property to get an invariant vector pricing; for instance, axiom (VP2) is not a convex condition. Therefore, the main step for ˜F will be to reduce the question to numerous cones of linear functionals using the following result.
@@
Let $G$ be a group with a representation on an ordered vector space $V$. Suppose that for every non-zero positive vector $v\in V$, the $G$-invariant ideal $V_{Gv}$ generated by the orbit $Gv$ admits a non-zero $G$-invariant positive linear functional.
@@ 1.4. Domains of definition
The critical reader might question two lifestyle choices that we made from the start. First, we defined conditional means $P(u|v)$ only when $u\leq v$, whereas the conditional probability $P(A|B)\in[0,1]$ is classically defined for any $A\subseteq X$, not only for $A\subseteq B$. In fact, Rényi’s original axioms [Rén55] are that $P(\cdot|B)$ is a probability measure on $X$ such that
@@
Moreover, $P(\boldsymbol{1}_{A}|\boldsymbol{1}_{B})=P(\boldsymbol{1}_{A\cap B}|\boldsymbol{1}_{B})$ holds for arbitrary $A,B\neq\varnothing$ when $V$ the space of step-functions on a set.
@@
Section˜2 sets the table with notation and background. Section˜3 introduces partial positive functionals and relates them to vector pricings and conditional means. A vectorial analogue of the Rényi order on measures is also introduced.
The main work starts with Section˜4, which discusses invariance. Section˜5 deals with stationarity. Section˜6 squares away questions of domains of definition.
@@ 2.1. Order
We also note that the dual of a representation is strictly speaking a representation of the opposite semigroup on the dual space, or equivalently a right action. We shall however use abusive language and speak of dual $G$-actions; for groups, this is of course the same as the usual action by the inverse element.
@@
Let $A\subseteq V$ be any subset. The ideal generated by $A$ is the intersection $V_{A}$ of all ideals of $V$ containing $A$. When $A=\{v\}$, the simpler notation $V_{v}$ is used. We record a few basic facts that will be used throughout.
@@
For (i), any ideal containing $A$ must contain $\operatorname{Span}(A)$ and hence also the right hand side. For the reverse inclusion, it suffices to observe that the right hand side is an ideal; it follows that it contains $V_{A}$.
@@
We call $V$ singular if there is $v\gneqq 0$ such that $J(v)=0$ for every positive linear functional $J$ on $V$. This does not happen in most familiar spaces (compare Section˜2.3 below), but there are extreme examples.
Example 2.2.
@@
2.2. The definition of vector pricings
@@ 3. Pricings vs means vs partial functionals
In this section, we establish dictionaries between three languages for the results of this article.
@@ 4. Invariance
We are now ready for the main results. For the record, we clarify the definition of invariance for vector pricings. Although we are mainly working with groups, we state this more generally for semigroups because this will be convenient for handling stationarity.
@@ 4.2. Invariant pricings
Let $G$ be a semigroup with a representation on an ordered vector space $V$. Suppose that for every non-zero positive vector $v\in V$, the $G$-invariant ideal $V_{Gv}$ generated by the orbit $Gv$ admits a non-zero $G$-invariant positive linear functional.
@@
We define a collection $\mathscr{C}$ of partial positive functionals as follows. Given any non-zero $v\geq 0$, let $V_{Gv}$ be the ideal generated by the orbit $Gv$; this is a $G$-invariant ideal. By assumption, there exists a non-zero $G$-invariant positive linear map $J\colon V_{Gv}\to\mathbf{R}$. We note that $J(v)\neq 0$; indeed otherwise by invariance $J(Gv)=\{0\}$ so that $J$ would vanish on the entire ideal $V_{Gv}$.
@@
Finally, the “local boundedness” condition postulates the existence of a non-zero element $x\in E^{+}$ whose orbit is bounded in the sense of topological vector spaces. Since the representation is order-bounded, there is $w\in V$ with $gv\leq w$ for all $g\in G$. Since the order is non-singular, we can find a positive linear functional $K\colon V\to\mathbf{R}$ with $K(v)\neq 0$. Now we claim that the restriction of $K$ to $V_{Gv}$ is the desired element $x\in E^{+}$. Indeed, weak-* boundedness of $Gx$ means that any basic neighbourhood of 0 of the form
@@
we must absorb $Gx$, i.e. $Gx\subseteq tU$ must hold for all sufficiently large $t>0$. Given such $v_{i}$ and $\epsilon$, we can find for each $i$ elements $g_{i,j}\in G$ with $j=1,\ldots,k_{i}$ and $\pm v_{i}\leq\sum_{j=1}^{k_{i}}g_{i,j}v$ because $v_{i}\in V_{Gv}$ (again ˜2.1(ii)). Now
@@ 5. Stationarity
We denote its orbits by $\langle\mu\rangle v=\{\mu^{*n}*v:n\in\mathbf{N}\}$ and observe that the orbit ideal $V_{\langle\mu\rangle v}$ is contained in the $G$-invariant ideal $V_{Gv}$ whenever $v\in V^{+}$. In particular, if the $G$-representation is order-bounded, then $\mu$ has order-bounded orbits. Moreover, if $\mu$ is non-degenerate, then we have $V_{\langle\mu\rangle v}=V_{Gv}$ because every $gv$ is in $V_{\mu^{*n}v}$ when $n$ is large enough relative to $g\in G$.
@@
In our current setting, we can deduce from this result a statement that is ready to be combined with Kesten’s criterion, as follows.
@@ 6.1. Step-functions
Given a non-empty subset $A\subseteq X$, we consider $P(\cdot|A)$ as a finitely additive probability measure on $A$, which has therefore a unique extension to a mean $P^{\prime}(\cdot|A)$ on the space $\ell^{\infty}(A)$. Thus in particular $P^{\prime}(\boldsymbol{1}_{B}|A)=P(B|A)$ when $B\subseteq A$.
@@ 6.2. Globalizing conditional means
$\forall u\in V^{+}:\kern 8.53581ptp_{v}(u)=\sup\{u\wedge nv:n\in\mathbf{N}\}\in V_{v}.$
@@
$P(p_{v}(u)|v)=P\big(\sup\{u\wedge nv:n\in\mathbf{N}\}\big|v\big)=\sup_{n\in\mathbf{N}}P\left(u\wedge nv|v\right),$
@@ 6.3. Staying positive
˜6.6 justifies the above claim because any natural extension of an invariant vector pricing provided by ˜E or ˜F for a non-trivial group would still be invariant and thus contradict ˜6.6.
@@
again absurd. All values of $r$ appearing in the above rather dim sum are $1$, then 0, so that no hypothetical $\infty-\infty$ could be used as a fig leaf. ∎